OpenStudy (thats_kyy):

A team t-shirt costs $3 per adult and $2 per child. On a certain day, the total number of adults (a) and children (c) who bought shirts was 100, and the total money collected was $275. Which of the following options represents the number of children and the number of adults who purchased team shirts that day, and the pair of equations that can be solved to find the numbers?

OpenStudy (mathstudent55):

You need two equations. One equation deals with the number of people, and the other equation deals with the cost.

OpenStudy (mathstudent55):

Let's start with the equation that deals with the number of people. There were a adults and c children How do you represent the sum of the numbers of adults and children using a and c?

OpenStudy (thats_kyy):

By Adding it @mathstudent55

OpenStudy (mathstudent55):

Yes, adding what?

OpenStudy (thats_kyy):

Umm Well I Really Dont Know But Ill Try By adding How many adults and children their are @mathstudent55

OpenStudy (mathstudent55):

Right. The number of adults is represented by the letter a. The number of children is represented by the letter c. The total number of people is the number of adults plus the number of children, so the total number of people is a + c

OpenStudy (mathstudent55):

Ok so far?

OpenStudy (thats_kyy):

Yea @mathstudent55

OpenStudy (mathstudent55):

We are told that the total number of people is 100. We know the total number of people is a + c, so we can write our first equation: a + c = 100

OpenStudy (thats_kyy):

Okay

OpenStudy (mathstudent55):

That was the equation that dealt with the number of people. Now we need an equation that deals with the costs.

OpenStudy (thats_kyy):

Okay

OpenStudy (mathstudent55):

1 adult shirt costs $3 * 1 2 adult shirts cost $3 * 2 3 adult shirts cost $3 * 3 a adult shirts cost $3 * a, or simply 3a

OpenStudy (thats_kyy):

Okay Im Getting It Now..

OpenStudy (mathstudent55):

Now we do the same for the children's shirts. 1 child shirt costs $2 * 1 2 child shirts cost $2 * 2 3 child shirts cost $2 * 3 a child shirts cost $2 * c, or simply 2c

OpenStudy (thats_kyy):

I Just Really Need To Know how Many Children And Adults They Are

OpenStudy (mathstudent55):

The cost of the adult shirts was 3a, and the cost of the children's shirts was 2c. The total cost is 3a + 2c We are told the total cost is $275, which gives us our second equation: 3a + 2c = 275

OpenStudy (thats_kyy):

okay

OpenStudy (mathstudent55):

Now we have a system of equations: a + c = 100 3a + 2c = 275

OpenStudy (thats_kyy):

Which Is a+c=100 3a + 2c = 275

OpenStudy (mathstudent55):

Now we solve this system of equations to find the number of adults and the number of children. We'll use the substitution method. Let's solve the first equation for a. a = 100 - c Now we substitute 100 - c in for a in the second equation: 3(100 - c) + 2c = 275 300 - 3c + 2c = 275 300 - c = 275 -c = -25 c = 25 Now we substitute c = 25 into the first equation: a + c = 100 a + 25 = 100 a = 75 Answer: 75 adults and 25 children

OpenStudy (thats_kyy):

OMG THANK YOUU THANK YOU SOO MUCH!!! But I Have a Couple More Im Stuck On Do You Mind Helping Me Please

OpenStudy (thats_kyy):

@mathstudent55

OpenStudy (mathstudent55):

You're welcome. I have another 15 minutes to help you. Please start a new post for a new problem. I'll go there to help you.